# Humans as a Calibration Pattern: Dynamic 3D Scene Reconstruction from Unsynchronized and Uncalibrated Videos

### v1

@@ 27                   Pop $\mathbf{G}_{l}[l]$ and add to $\mathbf{G}_{l}[k]$ after shift $\Delta{T}_{ij}$
We exhaustively find all pairwise alignments of human motions and store them into two $N\times N$ matrices. Each pairwise alignment employs a dynamic time warping (DTW) algorithm, which is widely used in speech recognition [35].
where time offset matrix $\Delta T\in\mathbb{Z}^{N\times N}$ stores the estimated relative time offsets and cost matrix $C\in\mathbb{R}^{N\times N}$ stores the cost of the alignment at the estimated time offset. To calculate the cost with the DTW algorithm, we use the distance between human states defined in Eq. 3. We assume that every video $V$ has the same frame rate, and we estimate time offset $\Delta{T}_{ij}$ as the most frequent warping time.
Then, we find the global sequence alignment of whole human motions with a greedy algorithm. First, we find an anchor index pair which has the minimum value $(i,j)=\operatorname*{argmin}_{(i,j)}C_{ij}$. Then, we incrementally add index pairs in increasing order with respect to the cost value.
$\Delta t=\text{Global Align}(C,\Delta T),$ (6)
where $\Delta t\in\mathbb{Z}^{N}$ is estimated time offsets that globally align human motions. A detailed pseudocode for the global alignment process is provided in Algorithm 1.
Camera pose estimation
3.3 Refinement with Dynamic NeRF Optimization
In this step, we further refine time offsets and camera poses, which are initialized properly in the initialization stage (Sec. 3.2) during the optimization of dynamic NeRF. We utilize K-Planes [8] as our dynamic NeRF representation. K-Planes uses six multi-resolution feature grids to represent 4D radiance fields. Namely, there are $\mathbf{P}^{xy}_{l}$, $\mathbf{P}^{yz}_{l}$, and $\mathbf{P}^{zx}_{l}$ for space-only grids, $\mathbf{P}^{xt}_{l}$, $\mathbf{P}^{yt}_{l}$, and $\mathbf{P}^{zt}_{l}$ for space-time grids, where $l\in[1,L]$ denotes the resolution level. For querying spatiotemporal point $\mathbf{q}=(\mathbf{x},t)$, we can obtain the features by
[…]
$f(\mathbf{q})=\bigoplus_{l=1}^{L}w_{l}f_{l}(\mathbf{q}),$ (10)
where $\mathcal{F}_{\sigma}$ and $\mathcal{F}_{\text{color}}$ are tiny MLP and $\gamma$ is a positional embedder introduced in original NeRF [28]. Then, we can obtain pixel value at time $t$ along a ray $\mathbf{r}_{t}^{i}=(\mathbf{o}_{t}^{i},\mathbf{d}_{t}^{i})$ with volume rendering formulation. $\mathbf{o}_{t}^{i}$ and $\mathbf{d}_{t}^{i}$ are camera center and ray direction which can be obtained with $i$th camera pose $R_{t}^{i},\tau_{t}^{i}$.
where $\mathbf{x}_{k}=\mathbf{o}_{t}^{i}+p_{k}\mathbf{d}_{t}^{i},p_{k}\in[\text{near},\text{far}]$ is the ray sample point on ray, $T_{k}=e^{-\sum_{j=1}^{k-1}\sigma_{j}\delta_{j}}$ is the accumulated transmittance, and $\delta$ is the distance between ray samples.
However, naïvely optimizing both dynamic NeRF and calibration parameters easily falls into bad local minima, although we have good initial points for camera poses and time offsets. We propose a progressive learning strategy for robust joint optimization of dynamic NeRF and calibration parameters.
where $\alpha=L(e^{\eta}-1)/(e-1)$, $\eta\in[0,1]$ is a normalized training step. This is in contrast to vanilla K-Planes [8] which uses all weights as 1. Our feature weighting strategy assigns big weights to low-frequency grids at the beginning of optimization stage and progressively increases weights of higher-resolution grids. Our strategy is inspired by the dynamic low-pass filter in BARF [21], but we instead modify the relative weights for grid resolution to effectively start from low-frequency features in multi-resolution K-Planes.
Furthermore, we propose a curriculum learning strategy for stable optimization. We first freeze camera poses and time offsets and only optimizes 4D NeRF for $s_{0}$ training steps. Then we add camera poses to the parameter list from $s_{0}$ iterations. Finally we add time offsets to learnable parameters and jointly optimize all parameters after another $s_{1}$ iterations. This progressive learning strategy is critical in preventing model from overfitting. This also aligns with observations from SCNeRF [11] that sequentially adds complexity to camera model.
4 Experiments
4.1 Experimental Setup
Dataset preparation
We use CMU Panoptic Studio dataset [14] which contains dynamic sequences of human movement captured from $31$ cameras surrounding the scene. Each of scenes contain various human motions, interactions with objects, and large portion of occlusions for some viewpoints. Note the view configuration is much harder than typical scenes such as NeRF blender scenes [28] that have $100$ train views on upper hemisphere or LLFF dataset [27]. We take subsequences from sports and office sequence and make new dataset containing five scenes Baseball, Tennis, Office1, Office2, and Office3. Each multi-view training video is $270$ frames long at $30$ FPS and starts from a random global timestamp to make unsynchronized setup. All video sequences are sampled to have at least 150 overlapping frames. Namely, maximum time offset between two videos is $120$ frames. The ground truth time offsets and camera poses are used for evaluation, and not provided to the system. We use $29$ or $30$ cameras for training dynamic NeRF, and one camera for test novel-view synthesis performance.
Implementation details
We undistort all training images before estimating human motion and training dynamic NeRFs with provided radial and tangential distortion parameters. We optimize K-Planes with Adam optimizer [7] with rendering loss in Eq. 14 and regularization losses including total variance in space, time smoothness loss, density L1 loss, and sparse transient loss introduced in original K-Planes [8]. We also use proposal sampling [2] to sample ray points for volume rendering. Although we first estimate camera poses for all timestamps in the initialization stage, we only use the camera poses at the first timestamp for simplicity since our dataset contains only static multi-view cameras. We further optimize one camera pose for each video at refinement. We jointly optimize calibration parameters and K-Planes for total 300k iterations and finish coarse-to-fine scheduling at 100k iterations. We use grid resolution of 240 for time axis, multi-scale level of $L=4$ or $5$ depending on scenes, at the maximum resolution of 384. We unfreeze camera pose parameters at $s_{0}=2$k steps, time offset parameters at $s_{0}+s_{1}=20$k steps.
4.2 Alignment Performance
Figure 5: We visualize estimated camera poses of the initialization step (left) and the refinement step (right) on Office2 scene. Blue and red frustums are estimation and ground truth, respectively.
Scene Rotation () Trans. (cm) $\Delta t$ (frames)
Init Refine Init Refine Data Init Refine
Baseball 5.302 0.328 22.582 0.158 58.10 0.700 0.025
Office1 3.883 0.420 19.754 0.171 68.35 3.448 0.031
Office2 8.889 0.660 37.971 0.372 56.42 1.300 0.029
Office3 3.946 0.363 19.242 0.165 60.03 0.800 0.026
Tennis 5.293 0.290 25.827 0.217 69.06 0.467 0.028
Average 5.463 0.412 25.075 0.217 62.39 1.343 0.028
Table 1: Quantitative results on camera pose estimation and time synchronization. We report the results from the first initialization stage (Init) and the second refinement stage (Refine).
First, we report the accuracy of time synchronization and camera pose estimation in Tab. 1. We measure errors after aligning estimated results to ground truth to recover original scale, orientation, and shifts. Both quantitative results in Tab. 1 (“Init” column) and visual illustration Fig. 5 (left) show that our initialization step can achieve reasonable first estimation. Our first alignment with human motion achieves camera pose calibration with a rotation error of less than $5.5^{\circ}$ and a translation error of less than $26$ cm on average, despite the absence of any input camera pose information. Furthermore, our initialization strategy can estimate fairly accurate time offsets (average 1.34 frames) even from severely unsynchronized videos (62.39 frames time offset on average). Starting from the initial estimation, our refinement with joint optimization of 4D NeRF can achieve near-perfect calibration as shown in Tab. 1 (“Refine” column) and Fig. 5 (right). Our approach aligns camera poses within rotation error of $0.4^{\circ}$ and translation error of $0.22$ cm on average. Also the error of estimated time offsets is less than 0.03 frames on average, which is equal to 1 millisecond.
As shown in Footnote 3 (first row), replacing the distance function degrades the time synchronization performance significantly considering we are using the same SMPL sequence.
We further investigate the robustness of our initialization stage. First, we measure time offset and camera pose estimation error after applying synthetic image degradation and color variations to each training video $V^{i}$. We apply gamma correction with randomly sampled within uniform range $\gamma\sim[0.35,2.1]$ for color variations and we add luminance noise and chromatic noise for image degradation. We also add noise to human shape parameter $\beta^{i}$ and pose parameter $\Theta_{t}^{i}$ to test robustness on accuracy of SMPL parameter estimation. Samples of degraded images and noised human motion can be found in Fig. 6.
Tennis scene Rotation () Trans. (cm) $\Delta t$ (frames)
Init Refine Init Refine Init Refine
L2 norm, $\beta,\Theta$ 5.382 0.656 26.130 0.261 2.100 0.030
SMPL noise, $\sigma=0.01$ 5.266 0.285 25.844 0.296 1.633 0.027
SMPL noise, $\sigma=0.02$ 5.307 1.240 25.825 0.276 0.533 0.028
SMPL noise, $\sigma=0.05$ 5.310 0.357 25.861 0.277 0.833 0.029
SMPL noise, $\sigma=0.1$ 5.241 0.518 25.383 0.257 2.100 0.030
SMPL noise, $\sigma=0.2$ 5.404 0.934 26.831 0.217 2.367 0.024
Degraded video 6.496 - 32.374 - 1.133 -
Default setup 5.293 0.290 25.827 0.217 0.467 0.028
Table 2: Quantitative results with input degradation [^3] We do not test refinement for the degraded video setup since NeRFs cannot fit scenes that have severe appearance inconsistency across videos., noised human motion, and different choice of human state distance.
We report quantitative results on robustness experiments in Footnote 3. Although the alignment performance with degraded videos is degraded compared to the original setup, still our approach produces reasonable time offsets and camera poses for initialization considering the severe appearance variation and quality degradation. This result supports our claim that human motion can serve as a robust mid-level representation for calibration. Our approach also shows comparable accuracy to the original data when the noise level is moderate ($\sigma\leq 0.05$). This results show that our aligning strategy is not sensitive to the quality of extracted human motion. Although time offset errors increase at severe noise level ($\sigma\geq 0.1$), our approach reduces the temporal misalignment drastically (from 69.06 of input data to less then 2.5). Furthermore, calibration parameters initialized from noised human motions were sufficiently accurate to recover precise time offsets and camera poses in the subsequent refinement step, as shown in Footnote 3.
Figure 7: Qualitative comparison of novel view synthesis performance.
Scene Sync-NeRF w/ GT pose Ours w/o prog. training Ours Oracle (GT pose & time)
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Baseball 21.17 0.833 0.189 19.23 0.577 0.412 27.20 0.919 0.072 26.80 0.925 0.065
4.3 Dynamic Novel-view Synthesis
To evaluate our dynamic scene reconstruction performance, we measure novel view synthesis errors. We measure photometric error for rendered images at test view only for timestamps that are overlapped by all of training videos since we take unsynchronized videos as input. We also conduct test-time optimization for accurate measurement that freezes NeRF parameters and optimizes only test camera poses and timestamps for small iterations before measure metrics. Since we are the first approach to reconstruct dynamic NeRF without any assumption of both camera poses and time synchronization, there is no existing work to compare with our method that has exactly same problem setup. Instead, we compare our approach to Sync-NeRF [17] which is a method that jointly optimizes time offsets with dynamic NeRF. Since Sync-NeRF cannot align camera poses, we relax our problem setup for Sync-NeRF by providing ground-truth camera poses. We also compare our method with the oracle, namely, K-Planes with ground-truth camera poses and time offsets provided by dataset.
Quantitative results of novel view synthesis in mean PSNR, SSIM [41], LPIPS [46] are reported in Tab. 3. Our approach achieves superior performance across all metrics compared to Sync-NeRF even if Sync-NeRF baseline uses ground-truth camera poses. It supports our claim that good initial point (time offset for Sync-NeRF comparison) is critical for gradient-based optimization. Without proper initialization of time offsets, Sync-NeRF cannot reconstruct high-quality dynamic scenes. Furthermore, our approach achieves almost on par performance to the oracle model. We observe that our approach achieves a slightly higher PSNR than the oracle in the Baseball scene. This is possible because the ground-truth camera poses in the real-world dataset are not perfectly accurate, and our method optimizes time offsets to subframe precision. We show qualitative results which are rendered at novel viewpoints in Fig. 7. Our approach achieves superior rendering quality compared to Sync-NeRF baseline and comparable results compared to the oracle model, which is aligned with the quantitative results. Sync-NeRF cannot calibrate time offsets without proper initialization and cannot reconstruct dynamic objects consequently. Videos rendered at novel viewpoints can be found in the supplementary material.
Additionally, we test our approach without our progressive learning strategy with initialized calibration parameters. Quantitative results and qualitative results are also included in Tab. 3 and Fig. 7, respectively. Even starting with proper camera poses and time offsets estimated by our initialization stage, without progressive training, the reconstructed dynamic scenes show inferior quality. This result validates the importance of the proposed progressive learning strategy.
5 Conclusion
We propose a practical solution to reconstruct neural dynamic 3D scenes containing humans from unsynchronized and uncalibrated multi-view videos. Given human motion of individual videos, we find the initial estimates of temporal offsets and camera poses, which are highly robust to occlusions or other adversaries. We then train 4D NeRF volume in a coarse-to-fine fashion which effectively stabilize the optimization process. During training, we progressively add the estimated parameters into a joint optimization routine and further refine them. We demonstrate that we can acquire reliable dynamic scene reconstruction in challenging setups performing on par with accurate calibration.
Limitations & future works
Although our initialization stage robustly estimates time offsets and camera poses for most cases thanks to the temporal aggregation, we cannot recover calibration parameters if SLAHMR [44] completely fails to recover human motion from video. Nevertheless, we anticipate that as advancements in human motion estimation continue, our approach will benefit from these improvements. Our formulation is general enough to be applicable even when all cameras are in motion; however, we have not yet empirically validated this scenario, and increased degrees of freedom may require additional optimization techniques. Additionally, integrating techniques that address varying appearance and occlusions (e.g., appearance embedding and uncertainty estimation [26]) could enable our method to reconstruct complex scenes from distributed videos, such as concert halls or sports games, directly from fan-captured internet videos.
Supplementary Material
Appendix A Implementation Details
In this section, we describe details that we could not address in the main manuscript. We describe additional details to train K-Planes, which is our dynamic NeRF representation in Sec. A.1, Procrustes alignment used in camera pose estimation in Sec. A.2, and details regarding evaluation in Sec. A.3.
A.1 Training K-Planes
We use $L=5$ spatial grid resolutions $[24,48,96,192,384]$ for Office1, Office2, Office3, and Tennis scenes, and we use $L=4$ grid resolutions $[48,96,192,384]$ for Baseball scene. We observe that Baseball scene converges well starting from the resolution of $48$. We use a single resolution, $240$, for the temporal grid in all scenes.
In addition to the weight scheduling described in the main manuscript, we also schedule the weights of regularization terms. We apply cosine scheduling that decreases weights to $1/100$ of its initial weights at the end of the scheduling. We use weights $0.01$ for distortion loss, $0.001$ for L1 loss in time planes, $0.001$ for total variance loss in spatial planes, $0.01$ for time smoothness loss, and $0.01$ for density L1 loss. We start scheduling of regularization from $100$k steps for Office1, Office2, Office3, and Tennis scenes and $50$k steps for Baseball scene, and end scheduling at $150$k steps.
A.2 Procrustes Alignment

### v2

@@ 3.2 Initialization Stage
We find all pairwise alignments of human motions and store them into two $N\times N$ matrices. Each pairwise alignment employs a dynamic time warping (DTW) algorithm, which is widely used in speech recognition [38].
where time offset matrix $\Delta T\in\mathbb{Z}^{N\times N}$ stores the estimated relative time offsets and cost matrix $C\in\mathbb{R}^{N\times N}$ stores the cost of the alignment at the estimated time offset. To calculate the cost with the DTW algorithm, we use the distance between human states defined in Eq. 2. We estimate time offset $\Delta{T}_{ij}$ as the most frequent warping time. If the framerate of two videos are different, we upsample 3D joint locations by linearly interpolating joint locations between observed frames of lower framerate videos.
Then, we find the global sequence alignment of whole human motions with a greedy algorithm. First, we find an anchor index pair which has the minimum value $(i,j)=\operatorname*{argmin}_{(i,j)}C_{ij}$. Then, we incrementally add index pairs in increasing order with respect to the cost value
$\Delta t=\text{{Global Align}}(C,\Delta T),$ (5)
where $\Delta t\in\mathbb{Z}^{N}$ is estimated time offsets that globally align human motions. A detailed pseudocode for the global alignment process is provided in the supplemental material.
Camera pose estimation
3.3 Refinement with Dynamic NeRF Optimization
In this step, we further refine time offsets and camera poses, which are initialized properly in the initialization stage (Sec. 3.2) during the optimization of dynamic NeRF. We utilize K-Planes [8] as our dynamic NeRF representation. K-Planes uses six multi-resolution feature grids for 4D NeRF. Namely, there are $\mathbf{P}^{xy}_{l}$, $\mathbf{P}^{yz}_{l}$, and $\mathbf{P}^{zx}_{l}$ for space-only grids, $\mathbf{P}^{xt}_{l}$, $\mathbf{P}^{yt}_{l}$, and $\mathbf{P}^{zt}_{l}$ for space-time grids, where $l\in[1,L]$ denotes the resolution level. We can obtain the features at querying spatiotemporal point $\mathbf{q}=(\mathbf{x},t)$ by
[…]
$f(\mathbf{q})=\bigoplus_{l=1}^{L}w_{l}f_{l}(\mathbf{q}),$ (9)
where $\mathcal{F}_{\sigma}$ and $\mathcal{F}_{\text{color}}$ are tiny MLPs and $\gamma$ is a positional encoder [29]. Then, we can obtain the pixel value at time $t$ along a ray $\mathbf{r}_{t}^{i}=(\mathbf{o}_{t}^{i},\mathbf{d}_{t}^{i})$ with volume rendering formulation. $\mathbf{o}_{t}^{i}$ and $\mathbf{d}_{t}^{i}$ are the camera center and ray direction, which can be obtained with $i$th camera pose $R_{t}^{i},\tau_{t}^{i}$
where $\mathbf{x}_{k}=\mathbf{o}_{t}^{i}+p_{k}\mathbf{d}_{t}^{i},p_{k}\in[\text{near},\text{far}]$ is the sample point on ray, $T_{k}=e^{-\sum_{j=1}^{k-1}\sigma_{j}\delta_{j}}$ is the accumulated transmittance, and $\delta$ is the distance between ray samples.
However, naïve joint optimization of dynamic NeRF and calibration parameters easily falls into bad local minima, although we have good initial points for camera poses and time offsets. We propose a progressive learning strategy for robust joint optimization.
where $\alpha=L(e^{\eta}-1)/(e-1)$, $\eta\in[0,1]$ is normalized training step. This is in contrast to vanilla K-Planes [8], which uses all weights as 1. Our feature weighting strategy assigns big weights to low-frequency grids at the beginning of the optimization stage and progressively increases weights of higher-resolution grids. Our strategy is inspired by the dynamic low-pass filter in BARF [21], but we instead modify the relative weights for grid resolution to effectively start from low-frequency features in multi-resolution K-Planes.
Furthermore, we propose a curriculum learning strategy for stable optimization. We first freeze camera poses and time offsets and only optimize 4D NeRF for $s_{0}$ training steps. Then, we add camera poses to the parameter list from $s_{0}$ iterations. Finally, we add time offsets to learnable parameters and jointly optimize all parameters after another $s_{1}$ iterations. This progressive learning strategy is critical in preventing the model from overfitting. This also aligns with observations from SCNeRF [11] that sequentially adds complexity to the camera model.
4 Experiments
4.1 Experimental Setup
Dataset
We evaluate our method on real-world datasets containing multi-view videos of dynamic scenes with human motions, including CMU Panoptic Studio [13], Mobile-Stage [46], and EgoBody [51]. We take subsequences of each multi-view video that start from random global timestamps, but all of them have overlapping frames. The ground truth time offsets and camera poses are used for evaluation and not provided to the system. CMU Panoptic Studio [13] contains various human motions, interactions with objects, and a large portion of occlusions for some viewpoints, which are captured from $31$ cameras located in the upper hemisphere surrounding the scene. We use $29$ or $30$ cameras for training dynamic NeRF and one camera for testing novel-view synthesis performance. Mobile-Stage dataset from 4K4D [46] contains three dancers captured by smartphone cameras with frontal and side views. We use $20$ cameras for training and one camera for testing. EgoBody [51] contains 4 or 5 static Kinect cameras and one moving head-mounted HoloLens2 camera, which has different intrinsics and framerate to Kinect. Since reconstructing NeRFs from sparse views (6 views) is challenging and an open problem, we focus on the evaluation of initialization stage performance on the EgoBody dataset. More details on datasets can be found in the supplemental material.
Implementation details
We optimize K-Planes with Adam optimizer [7] with rendering loss in Eq. 13 and regularization losses, including total variance in space, time smoothness loss, density L1 loss, and sparse transient loss introduced in original K-Planes [8]. We also use proposal sampling [2] to sample ray points for volume rendering. We jointly optimize calibration parameters and K-Planes for a total of 300k iterations and finish coarse-to-fine scheduling at 100k iterations. We use a grid resolution of 240 for the time axis and a multi-scale level of $L=4$ or $5$ depending on scenes at the maximum resolution of 384. We unfreeze camera pose parameters at $s_{0}=2$k steps, and time offset parameters at $s_{0}+s_{1}=20$k steps.
4.2 Alignment Performance
Figure 5: We visualize estimated camera poses of the initialization step (top) and after refinement (bottom) on various datasets. Blue and red frustums are estimation and ground truth, respectively.
Dataset Scene Rotation () Trans. $\Delta t$ (frames)
Init Refine Init Refine Data Init Refine
Baseball 5.30 0.328 22.6 cm 0.16 cm 29.04 0.700 0.025
Office1 3.88 0.420 19.7 cm 0.17 cm 33.25 3.448 0.031
Panoptic Office2 8.89 0.660 38.0 cm 0.37 cm 29.65 1.300 0.029
Studio Office3 3.95 0.363 19.2 cm 0.17 cm 32.93 0.800 0.026
Tennis 5.29 0.290 25.8 cm 0.22 cm 28.33 0.467 0.028
Average 5.46 0.412 25.1 cm 0.22 cm 30.64 1.343 0.028
Mobile-Stage Dance 5.65 1.707 0.053 0.002 24.76 0.455 0.214
EgoBody S22-S21-02 12.7 - 0.016 - 21.94 7.200 -
S32-S31-01 3.68 - 0.025 - 33.30 0.333 -
S32-S31-02 10.8 - 0.034 - 8.333 1.667 -
Average 9.07 - 0.025 - 21.19 3.067 -
Table 1: Quantitative results on camera pose estimation and time synchronization. We report the results from the first initialization stage (Init) and the second refinement stage (Refine). We set the size of scene bounding box as 1 to measure translation error for Mobile-Stage and EgoBody since the exact metric is not known.
First, we report the accuracy of time synchronization and camera pose estimation in Tab. 1. We measure errors after aligning estimated results to ground truth to recover original scale, orientation, and shifts. Both quantitative results in Tab. 1 (“Init” columns) and visual illustration in Fig. 5 (top) show that our initialization step can achieve reasonable first estimation across all datasets. Our initialization stage can robustly estimate calibration parameters regardless of the number of humans in the scene (one in Panoptic Studio, two in EgoBody, and three in Mobile-Stage), on various camera configurations (uniformly distributed in Panoptic Studio, sparse distributed in frontal and side views in Mobile-Stage, including moving camera in EgoBody). Our first alignment with human motion achieves camera pose calibration with a rotation error of less than $5.5^{\circ}$ and a translation error of less than $26$ cm on average in Panoptic Studio, despite the absence of any input camera pose information. Furthermore, our initialization strategy can estimate fairly accurate time offsets (average 1.34 frames) even from severely unsynchronized videos (average 30.64 frames offset). Also, our initialization stage consistently estimates accurate calibration parameters in Mobile-Stage and EgoBody dataset.
Starting from the initial estimation, our refinement with joint optimization of 4D NeRF can achieve near-perfect calibration as shown in Tab. 1 (“Refine” columns) and Fig. 5 (bottom). Our approach aligns camera poses within rotation error of $0.4^{\circ}$ and translation error of $0.22$ cm on average in Panoptic Studio. Also, the error of estimated time offsets is less than 0.03 frames on average, which is equal to 1 millisecond. Our method also shows highly accurate calibration after refinement in Mobile-Stage dataset as well.
As shown in Tab. 2 (first row), replacing the distance function degrades the time synchronization performance significantly considering we are using the same SMPL sequence.
We further investigate the robustness of our initialization stage. First, we evaluate our method after applying image degradation and color variations to each training video $V^{i}$. We apply gamma correction with randomly sampled within uniform range $\gamma\sim[0.35,2.1]$ for color variations and add luminance noise and chromatic noise for image degradation. We also add noise to human shape parameter $\beta^{i}$ and pose parameter $\Theta_{t}^{i}$ to test robustness on the accuracy of SMPL parameter estimation. Samples of degraded images and noised human motion can be found in Fig. 6. Additionally, we test mixed-framerate input using FPS=24 for five videos and FPS=30 for others.
Tennis scene Rotation () Trans. (cm) $\Delta t$ (frames)
Init Refine Init Refine Init Refine
L2 norm, $\beta,\Theta$ 5.382 0.656 26.130 0.261 2.100 0.030
SMPL noise, $\sigma=0.01$ 5.266 0.285 25.844 0.296 1.633 0.027
SMPL noise, $\sigma=0.02$ 5.307 1.240 25.825 0.276 0.533 0.028
SMPL noise, $\sigma=0.05$ 5.310 0.357 25.861 0.277 0.833 0.029
SMPL noise, $\sigma=0.1$ 5.241 0.518 25.383 0.257 2.100 0.030
SMPL noise, $\sigma=0.2$ 5.404 0.934 26.831 0.217 2.367 0.024
Degraded video 6.496 - 32.374 - 1.133 -
Mixed FPS 7.933 - 38.58 - 1.200 -
Default setup 5.293 0.290 25.827 0.217 0.467 0.028
Table 2: Quantitative results with different choice of human state distance, noised motion, input degradation, and mixed FPS setup.
Figure 7: Qualitative comparison on Mobile-Stage dataset
We report quantitative results on robustness experiments in Tab. 2. Although the alignment performance with degraded videos underperforms compared to the original setup, our approach still produces reasonable time offsets and camera poses for initialization, considering the severe appearance variation and quality degradation. This result supports our claim that human motion can serve as a robust mid-level representation for calibration. Our method also estimates reasonable calibration parameters from videos with mixed framerates. Our approach also shows comparable accuracy to the original data when the noise level is moderate ($\sigma\leq 0.05$). These results show that our aligning strategy is not sensitive to the quality of extracted human motion. Although time offset errors increase at severe noise levels ($\sigma\geq 0.1$), our approach drastically reduces temporal misalignment (from 28.33 of input data to less than 2.5). Furthermore, calibration parameters initialized from noised human motions were sufficiently accurate to recover precise time offsets and camera poses in the subsequent refinement step, as shown in Tab. 2.
Figure 8: Qualitative comparison of novel view synthesis performance on CMU Panoptic Studio dataset.
Dataset Scene Sync-NeRF w/ GT pose Ours w/o prog. training Ours Oracle (GT pose & time)
PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS PSNR SSIM LPIPS
Panoptic Studio Baseball 21.17 0.833 0.189 19.23 0.577 0.412 27.20 0.919 0.072 26.80 0.925 0.065
Mobile-Stage Dance 17.26 0.410 0.284 13.72 0.233 0.510 23.27 0.816 0.089 22.98 0.806 0.093
4.3 Dynamic Novel-view Synthesis
To evaluate our dynamic scene reconstruction performance, we measure novel view synthesis errors. We measure photometric error for rendered images at test view only for timestamps that are overlapped by all of the training videos since we take unsynchronized videos as input. We also conduct test-time optimization for accurate measurement that freezes NeRF parameters and optimizes only test camera poses and timestamps for small iterations before measuring metrics. Since we are the first approach to reconstruct dynamic NeRF without any assumption of both camera poses and time synchronization, there is no existing work to compare with our method that has exactly the same problem setup. Instead, we compare our approach to Sync-NeRF [16], which is a method that jointly optimizes time offsets with dynamic NeRF. Since Sync-NeRF cannot align camera poses, we relax our problem setup for Sync-NeRF by providing ground-truth camera poses. We also compare our method with the oracle, namely, K-Planes with ground-truth camera poses and time offsets provided by the dataset.
Quantitative results of novel view synthesis in mean PSNR, SSIM [43], and LPIPS [50] are reported in Tab. 3. Our approach achieves superior performance across all metrics compared to Sync-NeRF even if Sync-NeRF baseline uses ground-truth camera poses. It supports our claim that a good initial point (time offset for Sync-NeRF comparison) is critical for gradient-based optimization. Without proper initialization of time offsets, Sync-NeRF cannot reconstruct high-quality dynamic scenes. Furthermore, our approach achieves almost on par performance to the oracle model. Our approach achieves a slightly higher PSNR than the oracle in the Baseball scene in Panoptic Studio and Mobile-Stage. This is possible because the ground-truth camera poses in the real-world dataset are not perfectly accurate, and our method optimizes time offsets to subframe precision. We show qualitative results rendered at novel viewpoints in Figs. 8 and 7. Our approach achieves superior rendering quality compared to Sync-NeRF baseline and comparable results to the oracle model, which aligns with the quantitative results. Sync-NeRF cannot calibrate time offsets without proper initialization and cannot reconstruct dynamic objects consequently.
Additionally, we test our approach without our progressive learning strategy with initialized calibration parameters. Quantitative results and qualitative results are also included in Tab. 3 and Figs. 8 and 7, respectively. Even starting with proper camera poses and time offsets estimated by our initialization stage, without progressive training, the reconstructed dynamic scenes show inferior quality. This result validates the importance of the proposed progressive learning strategy.
5 Conclusion
We propose a practical solution to reconstruct neural dynamic 3D scenes containing humans from unsynchronized and uncalibrated multi-view videos. Given human motion extracted from individual videos, we find the initial estimates of temporal offsets and camera poses, which are highly robust to occlusions or other adversaries. We then train 4D NeRF volume in a coarse-to-fine fashion, which effectively stabilizes the optimization process. During training, we progressively add the estimated parameters into a joint optimization routine and further refine them. We demonstrate that we can acquire reliable dynamic scene reconstruction in challenging setups performing on par with accurate calibration.
Limitations & future works Although our initialization stage robustly estimates time offsets and camera poses for most cases, thanks to the temporal aggregation, we cannot recover calibration parameters if SLAHMR [48] completely fails to recover human motion from video. Nevertheless, we anticipate that as advancements in human motion estimation continue, our approach will benefit from these improvements. Our formulation is general enough to handle moving cameras; however, refining camera poses of moving cameras by joint optimization with 4D NeRF is left as future work. Additionally, integrating techniques that address varying appearances and occlusions (e.g., appearance embedding and uncertainty estimation [27]) could enable our method to reconstruct complex scenes from distributed videos, such as concert halls or sports games, directly from fan-captured internet videos. Also, extending our work beyond humans to generalize to other common objects (e.g., cars) would be also an exciting direction.
Supplementary Material
Appendix A Dataset Details
CMU Panoptic Studio
CMU Panoptic Studio dataset contains dynamic sequences of human movement from 31 cameras surrounding the scene. We take subsequences from sports and office sequence and make new dataset containing five scenes Baseball, Tennis, Office1, Office2, and Office3. There is one human in the scenes in Panoptic Studio dataset we used. Each multi-view training video is 270 frames long at 30 FPS and starts from a random global timestamp to make unsynchronized setup. All video sequences are sampled to have at least 150 overlapping frames. Namely, maximum time offset between two videos is 120 frames. We undistort all training images before estimating human motion and training dynamic NeRFs with provided radial and tangential distortion parameters.
Mobile-Stage
Mobile-Stage dataset from 4K4D contains three dancers captured by multiple smartphone cameras with frontal and side views. Some viewpoints have significant occlusion, and not all people are visible in certain viewpoints. The video lengths, FPS, and random global timestamp sampling strategy are identical to the setup in CMU Panoptic Studio dataset. We use 20 cameras for training and one camera for evaluation.
EgoBody
EgoBody dataset contains dynamic interaction between two humans that are captured from four (S22-S21-02) or five (S32-S31-01 and S32-S31-02) static Kinect cameras and one moving head-mounted HoloLens2 camera. HoloLens2 camera has different camera intrinsic from Kinect cameras and it has some missing frames. Also, there are extreme motion blurred frames in HoloLens2 camera videos. We estimate human pose for existing frame with SLAHMR and estimate human pose for missing frames by linear interpolation of 3D joint positions. Each multi-view training video is 200 frames long. They have at least 90 overlapping frames, in other words, maximum time offset between two videos is 110 frames.
Appendix B Implementation Details
B.1 Training K-Planes
We use $L$=5 spatial grid resolutions $[24,48,96,192,384]$ for Mobile-Stage dataset and Office1, Office2, Office3, and Tennis scenes of CMU Panoptic Studio dataset, and we use $L$=4 grid resolutions $[48,96,192,384]$ for Baseball scene of CMU Panoptic Studio dataset. We observe that Baseball scene converges well starting from the resolution of $48$. We use a single resolution, $240$, for the temporal grid in all scenes.
In addition to the weight scheduling described in the main manuscript, we also schedule the weights of regularization terms. We apply cosine scheduling that decreases weights to $1/100$ of its initial weights at the end of the scheduling. We use weights $0.01$ for distortion loss, $0.001$ for L1 loss in time planes, $0.001$ for total variance loss in spatial planes, $0.01$ for time smoothness loss, and $0.01$ for density L1 loss. We start scheduling of regularization from $100$k steps for Mobile-Stage dataset and Office1, Office2, Office3, and Tennis scenes of Panoptic Studio dataset, and $50$k steps for Baseball scene of Panoptic Studio dataset, and end scheduling at $150$k steps.
B.2 Global Alignment Algorithm
@@ 27                   Pop $\mathbf{G}_{l}[l]$ and add to $\mathbf{G}_{l}[k]$ after shift $\Delta{T}_{ij}$
We provide detailed pseudocode of the global sequence alignment of whole human motions for time offset estimation in Algorithm 1.
B.3 Procrustes Alignment